OpenStudy (anonymous):

sequence {X_n } is defined by X(n+1) = 2X(n)-X^2(n). If 00

OpenStudy (anonymous):

\[X _{n}\] is defined by \[X _{n+1}=2X _{n}-X ^{2}_{n}\] show that if \[0<X _{0}<1\] so is \[0<X _{n}<1\] for all integers n>0

OpenStudy (anonymous):

@IrishBoy123 wanna give me some hints with this one? i have tried using the same method..

OpenStudy (irishboy123):

try something like \(X_n = 2 - \frac{X_{n+1}}{X_{n}} \) as \(0<X_n <1\) then \(0< 2 - \frac{X_{n+1}}{X_{n}}<1\) do each inequality separately see if it leads somewhere...

OpenStudy (anonymous):

Something else you can try: Let \(f(x)=2x-x^2\). Then \(f'(x)=2-2x\). For \(0<x<1\), you have that \(f'(x)>0\), which would suggest that \(\{X_n\}\) is an increasing sequence (at least if \(0<X_n<1\)). This smells like an induction proof waiting to happen.